OpenStudy (anonymous):

Solve log3(x+4)-log3(x-2)=2log3 2

OpenStudy (anonymous):

How ?

hartnn (hartnn):

@onzuresjulie Hi :) \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\) do you know the log property, \(\log A-\log B=\log (A/B)\) ?

OpenStudy (campbell_st):

use log laws to simplify the problem \[\log_{3}(\frac{x + 4}{x -2}) = \log_{3} 2^2\] so you can then say \[\frac{x +4}{x -2} = 4\] solve the equation \[x + 4 = 4(x -2)\]

OpenStudy (anonymous):

I got -2.4

OpenStudy (campbell_st):

distribute the right hand side x + 4 = 4x - 8 add 8 to both sides x + 12 = 4x can you solve this..?

OpenStudy (anonymous):

Wow! Thank you!!:D

OpenStudy (anonymous):

X=4